Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Sum of a fair die

Expert replies
by XLogic » Mon Apr 16, 2012 7:41 pm
A fair die is rolled twice. What is the probabilty that the sum of the landed sides are divisible by 2.

A) 6
B) 1/36
C) 1/6
D) 1/192
E) 36

OA:[spoiler]Really?[/spoiler]
my post helped --> thank me!
don't thank me --> my post = what the..??
Join the discussion
Source: — Problem Solving |

by Bill@VeritasPrep » Mon Apr 16, 2012 8:25 pm
With two rolls, we have four possible outcomes:

Odd-Odd
Even-Even
Odd-Even
Even-Odd

Each of these has a probability of 1/4:

Odd-Odd = 1/2 * 1/2 = 1/4
Even-Even = 1/2 * 1/2 = 1/4
Odd-Even = 1/2 * 1/2 = 1/4
Even-Odd = 1/2 * 1/2 = 1/4

Which of them give us an even sum? Odd-Odd (1 + 3 = 4) and Even-Even ( 2 + 4 = 6) do, but Even-Odd and Odd-Even (1 + 2 = 3) do not. Thus, the probability of an even sum should be 1/4 + 1/4 = 1/2, which is not an answer choice. Is it a six-sided die numbered 1-6?
Join Veritas Prep’s 2010 Instructor of the Year, Matt Douglas for GMATT Mondays

Visit the Veritas Prep Blog

Try the FREE Veritas Prep Practice Test
Join the discussion

by Anurag@Gurome » Mon Apr 16, 2012 8:28 pm
XLogic wrote:A fair die is rolled twice. What is the probabilty that the sum of the landed sides are divisible by 2.

A) 6
B) 1/36
C) 1/6
D) 1/192
E) 36

OA:[spoiler]Really?[/spoiler]
The sum of the landed sides can be divisible by 2 for the following combinations:
(1, 1), (1, 3), (1, 5), (3, 1), (5, 1)
(2, 2), (2, 4), (2, 6), (4, 2), (6, 2)
(3, 3), (3, 5), (5, 3)
(4, 4), (4, 6), (6, 4)
(5, 5)
(6, 6)

Therefore, probabilty that the sum of the landed sides are divisible by 2 = 18/36 = [spoiler]1/2[/spoiler]
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by XLogic » Mon Apr 16, 2012 10:31 pm
My mistake. The question is incomplete.

I think Anurag has it right.

@Anurag, is there a way to do this without listing everything out like that? i.e., using permutations + probability? Thanks.
my post helped --> thank me!
don't thank me --> my post = what the..??
Join the discussion

by Bill@VeritasPrep » Tue Apr 17, 2012 7:21 am
My method didn't require writing anything out, just figuring out the possible arrangements and the probabilities for a given roll.
Join Veritas Prep’s 2010 Instructor of the Year, Matt Douglas for GMATT Mondays

Visit the Veritas Prep Blog

Try the FREE Veritas Prep Practice Test
Join the discussion

by aneesh.kg » Tue Apr 17, 2012 7:24 am
The simplest method of all:

In order to get a sum of even, you should either have even on first dice & even on second dice OR odd on first dice & odd on second dice.

P(Sum is even) = P(even on 1st)*P(even on 2nd) + P(odd on 1st)*P(odd on 2nd)
= (3/6)*(3/6) + (3/6)*(3/6)
= 1/2*1/2 + 1/2*1/2
= 1/4 + 1/4
= 1/2
Aneesh Bangia
GMAT Math Coach
[email protected]

GMATPad:
Facebook Page: https://www.facebook.com/GMATPad
Join the discussion

by aneesh.kg » Tue Apr 17, 2012 7:45 am
Or, even intuitively, if you have equal number of even and odd numbers on a Dice, the Sum should be even in exactly half of the cases. So, the probability should be 50% or 1/2.

The intuition can be explained by the knowledge of
Sum being even in a (even X even) or (odd X odd) case,
and
Sum being odd in a (odd X even) or (even X odd) case.

So, out of four equally possible cases, 2 favour us. Therefore, Probability = 1/2.

Anyway, if this doesn't make too much sense, follow the method given above.
Aneesh Bangia
GMAT Math Coach
[email protected]

GMATPad:
Facebook Page: https://www.facebook.com/GMATPad
Join the discussion

by ronnie1985 » Tue Apr 17, 2012 8:54 am
The sum of numbers from 1 to 6 taken 2 at a time is either even or odd, probability s 1/2 which is not in the option. Something is wrong with the question.
Follow your passion, Success as perceived by others shall follow you
Join the discussion